Zn reacts with hydrochloric acid.
Zn(s) + 2 HCl(aq) → ZnCl 2(aq) + H 2(g)
What volume of 1.53 M HCl(aq) will react with 15.0 g Zn(s)?
Answer #1

Moles of Zn = mass/molar mass = 15/65.38 = 0.2294

From reaction;

1 mole Zn requires 2 moles of HCl

So, 0.2294 moles of Zn will require = 2*0.2294 = 0.4588 moles of HCl

Molarity of HCl = moles/volume in liters

1.53 = 0.4588/V

V = 0.2999 liters = 0.3 liters = 300 ml (approx) ....Answer